OpenStudy (anonymous):

solve for x: log(x+4)-log(x+3)=2

OpenStudy (anonymous):

log(x+4)=log(x)+log4 log(x+3)=log(x)+log3 so = log(x)+log(4) + log(x)+log(3) =2log(x)+(log(4)+log(3)) 2log(x)=-log(4)-log(3) log(x)=(-log(4)-log(3))/2 x=10 ^-(log(4)-log(3))/2

OpenStudy (anonymous):

\[x=10^{-(\log4+\log3)/2}\]

OpenStudy (anonymous):

that is what I got